OpenStudy (eva12):

1)Which of the following transformations are linear? A. {y1=6x1,y2=5 B. {y1=2x1+x2,y2=2 C.{y1=2,y2=7,y3=9 D. {y1=x1+10,y2=x2 E. y1=−3x,y2=14x1,y3=8x1 F.y1=5x1−4x2+3x3,y2=10x2−2x3,y3=−9x1−7x2

OpenStudy (eva12):

are b,e,f are linear?

OpenStudy (loser66):

To test whether it is a linear transformation or not, by definition, we need to check \[T(\alpha u + \beta v )= T(\alpha u)+ T(\beta v)\] or give out a counter example for a) let u =(1,2) v = (0,1) so, u+ v= (1,3) T(u+v) = (6,5) \(\neq\)T(u) + T(v) = (6,5)+(0,5) = (6,10) --> T is not a linear transformation for b) let u =(1,2) v= (0,1) u+ v = (1,3) T(u+v) = (2*1+3, 2)= (5,2)\(\neq\) T(u) + T(v) = (2*1,2)+(2*0+1)= (2,3) for c) it is not a linear transformation and you have to prove it step by step let u = (x1, x2,x3) v = (t1, t2,t3) \[T( u + v) = (2,7,9) \neq T( u) + T ( v)= ....\]

OpenStudy (loser66):

I may exagerate the problem, but I failed the course because of this stuff so that I am so careful when solving it.

OpenStudy (dan815):

I see so linearity jus means dos it follow principle of superposition

OpenStudy (eva12):

then if i use A.{y1=6x1,y2=5 from the question T(y1,y2)=(6x1,5) do i plug in random numbers then use T(u+v)=T(u)+T(v) ?